# Further Comments on Yablo’s Construction ††thanks: File: WAB, [Sch25], arXiv 2504.10370

### v3

@@ Abstract
1 Overview
These notes are based on [Sch22] and [Sch23b], the reader may have copies of both ready for more examples and discussions. All these papers are, of course, based on Yablo’s seminal [Yab82].
See also the author’s earlier versions on arXiv.
In Section 2 (page 2), we first present the background, its logical and graph theoretical side, and their interplay. We also discuss a useful third truth value, which expresses that neither nor is consistent.
@@ 1 Overview
In Section 3 (page 3), we discuss elementary contradiction (cells). Essentially, these are variations of Yablo’s triangles. The simplest contradictions are too simple, they have an “escape possibility”, i.e., combining them will leave one possibility without contradictions. Diamonds, constructions with four sides, will not work because they need a “synchronization”, which is not available in our framework.
In Section 4 (page 4), we discuss how to put contradiction cells together to obtain graphs, which code the liar paradox. Basically, we follow Yablo’s construction, but give a finer analysis, discuss its properties, and again escape possibilities (in indirect composition).
We may start the construction with some simple but we will automatically end in a Yablo like construction, when satisfying the basic requirements (C1) and (C2) (see below, Section 1.1 (page 1.1)). This is discussed in the sections on “Saw Blades”, see Section 7.6 in [Sch22] and Section 3 in [Sch23b].
@@
1.1 Basic Terminology
@@
We denote by (C1) (for the condition that has to be contradictory, and by (C2) the condition that has to be contradictory, which means by that all have to be contradictory, roughly, in graph language, that all paths from have to lead to a contradiction. (More precisely, see Section 4.1 (page 4.1), (3).)
1.2 Comments On the Strategy
These comments will be clearer in hindsight, after having read the details. Still the best place to put them seems to be in the introduction.
The comments on the construction in the limit solve a seeming contradiction by chosing the right level of abstraction, i.e. by arguing on the level of instances. We do not show that there cannot be any other way to do this - we do not know, we just see no other possibility.
1.2.1 Expressing “x is false” In an Acyclic Graph
is sometimes expressed by and by a contradiction.
How can we express that is wrong in a finite acyclic graph?
Consider the set of minimal elements below , and the formula corresponding to the graph in DNF of these elements. This might be
(the graph is trivial, just
or a tautology or contradiction.
So the only way to express that a is wrong is by a contradiction.
We want to express by an infinite graph that both and are inconsistent.
By (102.2), this must be expressed by suitable contradictions. Therefore, we treat such contradictions as basic building blocks, as Yablo did.
The problems are then
- (1)
This view is too abstract, we have to look at details, as they are given in the proof of validity for Yablo’s construction, or our reconstruction in Section 4 (page 4).
We have to look at instances.
- (2.1)
- (2.2)
- (2.4)
2 Background: Graph, Logik
Most of the material in this section is either common knowledge, or was already covered in [Sch23b].
2.1 The Logical Side
@@
We will see that we need infinite formulas (of infinite depth) to have nodes to which we cannot attribute truth values. See Fact 2.4 (page 2.4).
@@
2.2 The Graph Side
Definition 2.1
Nodes stand for propositional variables.
Usually, the are (possibly infinite) conjunctions of propositional variables or (in most cases) their negations, which we write etc. We often indicate the negated variables in the graph with negated arrows, like etc. Thus, usually stands for
@@ 2.2 The Graph Side
in above definition cannot be replaced by etc, as this example shows (we argue semantically, with the central conflictual construction in Yablo’s paper, Yablo Cell, see Section 1.1 (page 1.1), (2)):
Example 2.2
Consider the basic construction of a contradiction (used by Yablo and here, see Section 1.1 (page 1.1).
stands for so which is impossible.
If we negate then so is possible.
From the graph perspective, we have two paths in from to and
We add now to so thus so and and is not contradictory any more.
Definition 2.2
Formally:
Let be paths as usual.
If then if then
Let be the concatenation of and Then iff and - otherwise.
If all arrows are negative, then iff the length of is even.
Definition 2.3
Remark 2.2
We saw already in Example 2.2 (page 2.2) that it is not sufficient for a “real” contradiction to have two contradictory paths.
We need
Notation 2.1
Fact 2.3
(Simplified).
Consider three paths, for simplicity with same origin, i.e.
No contradiction loops of length 3.
(Trivial, we have only 2 truth values).
Example 2.3
(Example 7.4.2 in [Sch22].)
Let
Obviously, this may be generalized to paths.
Consider Yablo’s original construction:
Example 2.4
2.3 Interplay of the Graph and the Logical Side
We need infinite depth and width in our constructions:
Fact 2.4
The construction needs infinite depth,
the logic as used in Yablo’s construction is not compact,
Proof: See Fact 1.4, 858 in [Sch23b].
2.4 A Third Truth Value
Remark 2.5
We consider a third truth value, in addition to and
Let be the truth value of
Let be an abbreviation for and
Let
We have
analogously.
Likewise
We thus have:
and
Using the operations for truth values
and above results, we have
(or and thus the order
(and finally the inverse operations
Definition 2.4
We loosely use the expression “escape possibility”, “escape path” etc. to indicate a construction which avoids a desired property through a (usually infinite) sequence of decisions.
Examples:
A simple example is the trivial contradiction, so (See Definition 3.1 (page 3.1).) If we continue with so etc. we have an escape sequence etc., which never meets a contradiction.
“Pipelines”: We may try to avoid infinite branching by one branch used to branch infinite often, but sequentially.
Say, instead of a disjunction of infinitely many arrows we construct one arrow and branch or or etc., but now the sequence …. avoids hitting any
Remark 2.6
Suppose we have an escape problem, like thus
This, however, is not possible.
3 Elementary Cells
3.1 General Remarks
We try to describe here the basic constructions of contradictions.
As said in Fact 2.1 (page 2.1) (and re-written graph-theoretically), these consist of two branches with common origin, which meet again, and have different polarity. We call such constructions cells.
Second, we will define a simple hierarchy of such cells, and will allow a cell only the use of simpler cells as substructure. This prevents “cheating”. (See Remark 3.1 (page 3.1).) Thus, we look at those cells only which allow to construct a Yablo-like structure without the use of more complicate cells.
Definition 3.1
We will argue that it suffices to consider these types of cells, see Remark 3.3 (page 3.3) and Section 3.2 (page 3.2). Furthermore, we can also neglect all but the (slightly generalized) Yablo Cells, as we will see in the present section.
Remark 3.1
Suppose we add to the Yablo construction an additional node (or will have the same properties (i.e. truth value as but this addition will not contribute anything to the construction.
To prevent this, we require that it is possible to construct a Yablo structure using only the contradiction cell discussed (and perhaps simpler contradiction cells).
Remark 3.2
We recall again the requirement (C2), has to be contradictory. As (or and thus (or we add the requirement that and (or and are each contradictory.
Thus, Yablo cells and diamonds are different from the simple contradiction.
Yablo cells and diamonds are different from each other as long as we require in the latter case that the truth value of too is and not just False. We will see that making and leads to problems.
(Conversely, making in a Yablo cell not contradictory makes the construction, of course, equivalent to the simple contradiction.)
Remark 3.3
False, True
We present a systematic treatment of variants of the Yablo triangle (and the diamond).
In the Yablo construction, we attach to the Yablo triangle to and constructions, which make We also consider here the cases where we end by or and set e.g. etc., see (2.3.1) below.
We consider:
Do we have a contradiction for
Do we have contradictions for
Do we have escape possibilities?
Do we obtain
- (2.2.5)
We consider the cases in all possible combinations:
Case is equivalent to case
Thus, we look at e.g. case i.e. and etc., and consider whether the requirements in (2.2) above are satisfied.
We first look only at the minimal requirements
has to be contradictory, condition (2.2.1) above.
must not be classical condition (2.2.4) above.
The cases:
Consider so for any so these cases are classical.
Consider For is classical, so is, too. so Thus, is the only non-classical case among
Consider In cases so too. In case so In case so So is the only classical case among the
So the interesting cases left are
We now check for local contradiction in the case
This is simpler, we just have to consider the cases (e.g.) the are omitted, as we look for local contradictions, is anything, and, of course,
The case is known (Yablo construction, all cases except and
Case and we have no contradiction.
Case and we have a contradiction.
So are contradictory.
Next, we have to look at and see if every path from leads to a contradiction, condition (2.2.2) above.
We have to check it must not be
So and will not work, i.e.
Consider i.e. Then in and so this does not work. so this works. so this works.
Consider i.e. Then or and all cases work, except
In summary: in exactly the following cases:
We summarize the three conditions above:
(2.2.4): are ok.
(2.2.2): are not ok, so we learn nothing new beyond (2.2.4).
(2.2.1): are ok, so only is eliminated from those satisfying (2.2.4).
In Summary, only satisfy all three conditions, (2.2.1), (2.2.2), (2.2.4).
We look at these cases.
Here, and in (2.3.6), stands for which is of type likewise as we will append in the full structure at and a construction with value
In all cases
Thus,
Thus,
Thus,
identical to
Thus,
Thus,
We look at the Diamond.
Thus,
The modified Diamond is the same, only so and so we have a Yablo cell.
Summary of the (in this context) important constructions
3.2 Comments on Yablo Cells
Prerequisites:
We consider a triangle , , but leave open if the arrows are positive or negative. We require that is of type and that the triangle is contradictory for the case as well as for both paths und have to lead to a contradiction (conditions (C1) and (C2)).
3.2.1 Branching points
We consider the hierarchy:
intermediate point - branching point - branching point with contradiction
If we have a contradiction on only at the situation is again equivalent to the trivial contradiction.
More general, if and contradict each other, and both are for x- valid paths (i.e. no contradiction at then we have an escape possibility.
The points
Suppose we branch at so we have, in addition to the triangle, some
Case 1: If then AND, thus if x- then OR:
Consider , , so, whatever the choice in we have what we need.
Without a contradiction at we have an escape possibility for (see above), e.g. continue the same way: , ,
Case 2: implies OR: (Yablo)
This however, is possible only a finite number of times, e.g. then etc., like “Pipeline”, Example 7.37, and Section 7.6.3 in [Sch22] or Example 1.5 and Section 3.4 in [Sch23b]. See also Definition 2.4 (page 2.4) above. This constructs an infinite sequence of choices which offers an escape possibility.
Thus we need infinite branching (and non-classical logic).
3.2.2 Paths
- (1)
We branch on , and have , ,
Case 1: OR:
We make a copy of the structure for
Case 2: AND:
So x- OR:
We make a copy of the structure for
Consider e.g. , ,
Diagram 3.1
3.2.3 All Sides Of the Triangle Have to Be Negative
Consider , . Suppose all are of the type
To obtain a contradiction for the triangle needs one or three negative sides.
Assume it has only one negative side.
Let , be a contradiction for (condition (C1)) and , be a contradiction for (for condition (C2)).
Suppose is positive, so we have or and, analogously, or (they have to be of the same type).
We leave aside the condition between and it is only the branching at we are interested in.
Thus, we have or
In the first case so and both and are possibilities, and, as in the simplest contradiction for we have an escape situation (see Definition 2.4 (page 2.4), (1)) (” watered down “by the additional possibilities and
In the second case so and the situation is analogous to the first case.
Suppose is negative, so we have and analogously
Now, we have to contradict for the case but for this we need a third negative arrow so this is impossible, and leads back to the usual Yablo cell.
There is a way out. Instead of considering we continue by This is equivalent to the Yablo cell by
But this is “cheating”.
3.2.4 The Problem with Diamonds
Remark 3.4
The simplified version with recursion only on the left is nothing but the Yablo triangle:
Diagram 3.2
Nested Diamonds, Details
Diagram 3.3
4 Combining Cells
4.1 Overview
We follow Yablo’s construction.
(C1): We have to make contradictory.
(This is not an important simplification, as more complicated constructions will finally use our simple one, too. See e.g. Diagram 4.3 (page 4.3) and its discussion.)
As Yablo did, we set and
(C2) We have to make contradictory.
Probabilistically, of the possibilities at are open.
Etc.
See below and Section 3.3.3. in [Sch23b].
The right level of abstraction:
The author tried first to work towards a representation result using single arrows. However, the following result suggests that one should work with (perhaps composite) paths instead of arrows: a single negative arrow may be replaced by the following diagram:
So
Thus, we may replace negative arrows recursively by arbitrarily deep constructions using negative arrows again, complicating the same construction arbitrarily.
The same is true for all other logical operators.
(See Section 4 in Sch23b.)
4.2 In More Detail
We now consider the inductive construction of the Yablo structure, see Diagram 4.1 (page 4.1) and Diagram 4.2 (page 4.2)
abbreviated
has to lead to a contradiction for
preparation,
finish,
(branch has to be contradicted for add
has to lead to a contradiction for
preparation,
finish,
(branch has to be contradicted for add
(branch has to be contradicted for add
Construction 4.1
Summary:
Every is head of a Yablo cell with knee
The construction has infinite depth and branching.
Diagram 4.1
Diagram Inductive Construction 1
Diagram 4.2
Diagram Inductive Construction 2
4.2.1 Composition
The prerequisites for the constructions in the diagrams Diagram 4.1 (page 4.1) etc. are negative Paths, und formulas.
These prerequisites hold also in our generalization to paths (modulo some simplifications for origins at etc.), so these constructions are also valid in the generalization to paths.
Remark 4.1
The condition “Existence of negative paths” is not trivial. Suppose that and are both negative, and is the only path from to then the condition is obviously false. The author does not know how to characterize graphs which satisfy the condition. Some kind of “richness” for the graph will probably have to hold. (This is covered in Yablo’s construction by transitivity.)
Example 4.1
Diagram 4.3
Diagram Postponed Contradictions
4.3 Beyond Conjunctions
We work with disjunctive normal forms.
@@ 4.3 Beyond Conjunctions
5 Conjecture

### v4

@@ Abstract
1 Introduction
Please note that arXiv pdf files sometimes have problems with some of the diagrams. In that case, please consult the html version.
1.1 Outline
These notes are based on [Sch22] and [Sch23b], the reader may have copies of both ready for more examples and discussions. See also the author’s earlier versions on arXiv. In particular, the reader is referred there for a discussion of more complicated formulas attached to the nodes. We liberally use material from these sources.
All these papers are, of course, based on Yablo’s seminal [Yab82].
In Section 1.2 (page 1.2), we first present the background, its logical and graph theoretical side, and their interplay. We also discuss a useful third truth value, which expresses that neither nor is consistent.
@@ 1.1 Outline
In Section 2 (page 2), we discuss elementary contradictions (cells). Essentially, these are variations of Yablo’s triangles. The simplest contradictions are too simple, they have an “escape possibility”, i.e., combining them will leave one possibility without contradictions. Diamonds, constructions with four sides, will not work because they need a “synchronization”, which is not available in our framework.
In Section 3 (page 3), we discuss how to put contradiction cells together to obtain graphs, which code the liar paradox. Basically, we follow Yablo’s construction.
We then generalize slightly, see Remark 3.1 (page 3.1), but show that this general construction contains Yablo’s construction, so it is minimal. (See also the discussion of “Saw Blades” in Section 7.6 of [Sch22] and Section 3 in [Sch23b].)
In both cases, we show that - within the chosen variant - the construction is necassary.
@@
There is an important property for the graphs, which we need, probably a sort of “richness”, see Remark 1.5 (page 1.5). This property holds trivially in Yablo’s construction by transitivity. In our approach, it will hold by the construction of a suitable injection, see Section 4 (page 4), Condition 4.1 (page 4.1) (2).
1.2 Basic Definitions and Facts
1.2.1 Terminology
@@
We denote by (C1) (for the condition that has to be contradictory, and by (C2) the condition that has to be contradictory, which means by that all have to be contradictory, roughly, in graph language, that all paths from have to lead to a contradiction. (More precisely, see Section 3.1 (page 3.1), (3).)
1.2.2 The Logical Side
@@
We will see that we need infinite formulas (of infinite depth) to have nodes to which we cannot attribute truth values. See Fact 1.6 (page 1.6).
Notation 1.1
@@
Fact 1.2
Let and we want to show that is contradictory, then it suffices to show that a subset of the is inconsistent - this corresponds to the fact that (X any truth value.)
Let and we want to show that is contradictory, then we have to show that ALL are contradictory. Again, this corresponds to
The following example illustrates the propagation of required contradictions.
@@
1.2.3 The Graph Side
Definition 1.1
Nodes stand for propositional variables.
Usually, the are (possibly infinite) conjunctions of propositional variables or (in most cases) their negations, which we write etc. We often indicate the negated variables in the graph with negated arrows, like etc. Thus, usually stands for more precisely.
Example 1.2
in above definition cannot be replaced by etc, as this example shows (we argue semantically, with the central conflictual construction in Yablo’s paper, Yablo Cell, see Section 1.2.1 (page 1.2.1), (2)):
Example 1.3
Consider the basic construction of a contradiction (used by Yablo and here, defined “officially” in Definition 2.1 (page 2.1).
stands for so which is impossible.
If we negate then so is possible.
From the graph perspective, we have two paths in from to and
We add now to so thus so and and is not contradictory any more.
Definition 1.2
Formally:
Let be paths as usual.
If then if then
Let be the concatenation of and Then iff and - otherwise.
If all arrows are negative, then iff the length of is even.
Definition 1.3
Remark 1.3
We saw already in Example 1.3 (page 1.3) that it is not sufficient for a “real” contradiction to have two contradictory paths.
We need
Fact 1.4
(Simplified).
Consider three paths, for simplicity with same origin, i.e.
No contradiction loops of length 3.
(Trivial, we have only 2 truth values).
Example 1.4
(Example 7.4.2 in [Sch22].)
Let
Obviously, this may be generalized to paths.
Consider Yablo’s original construction:
Example 1.5
Example 1.6
This Example shows that infinitely many finitely branching points cannot always replace infinite branching - there is an infinite “procrastination branch” or “escape branch”. This modification of the Yablo structure has one acceptable valuation for
Let as usual, and introduce new
Let with
If then by so, generally,
if then and
If then so if etc., so, generally,
if then or
Suppose now then for all and for all By there is and a contradiction, or for all again a contradiction.
But is possible, by setting and for all
Thus, replacing infinite branching by an infinite number of finite branching does not work for the Yablo construction, as we can always chose the “procrastinating” branch.
Diagram 1.1
Escape Path
Remark 1.5
For a representation result, we need enough paths in the graph considered to form contradictions. This is not trivial. Suppose that and are both negative, and is the only path from to then the condition is obviously false. The author does not know how to characterize graphs which satisfy the condition directly. Some kind of “richness” for the graph will probably have to hold. (This is covered in Yablo’s construction by transitivity.)
1.2.4 Interplay of the Graph and the Logical Side
We need infinite depth and width in our constructions:
Fact 1.6
The construction needs infinite depth,
the logic as used in Yablo’s construction is not compact,
Proof
Let be a minimal element, then we can chose an arbitrary truth value and propagate this value upwards. If there are no infinite descending chains, we can do this for the whole construction.
The logic as used in Yablo’s construction is not compact:
Trivial. (Take This is obviously inconsistent, but no finite subset is.).
It is impossible to construct a Yablo-like structure with classical logic:
Take an acyclic graph, and interpret it as in Yablo’s construction. Wlog., we may assume the graph is connected. Suppose it shows that cannot be given a truth value. Then the set of formulas showing this does not have a model, so it is inconsistent. If the formulas were classical, it would have a finite, inconsistent subset, Define the depth of a formula as the shortest path from to this formula. There is a (finite) such that all formulas in have depth Give all formulas of depth (arbitrary) truth values, and work upwards using truth functions. As the graph is acyclic, this is possible. Finally, has a truth value.
Thus, we need the infinite
The following example illustrates (partially) unfinished contradictions, and resulting escape possibilities.
Example 1.7
Diagram 1.2
Diagram Postponed Contradictions
1.2.5 A Third Truth Value
Remark 1.7
We consider a third truth value, in addition to and
Let be the truth value of
Let be an abbreviation for and
Let
We have
analogously.
Likewise
We thus have:
and
Using the operations for truth values
and above results, we have
(or and thus the order
(and finally the inverse operations
1.3 Strategy
These comments will be clearer in hindsight, after having read the details. Still the best place to put them seems to be in general remarks, and go into details later on.
The comments on the construction in the limit solve a seeming contradiction by chosing the right level of abstraction, i.e. by arguing on the level of instances.
1.3.1 Expressing “x is false” In an Acyclic Graph
In an acyclic graph, we cannot express something like we have to use different means to express that is impossible.
As we can always fill in truth values (including from the bottom, we cannot force by looking upwards of (or sideways, i.e. there is no monotone path connecting with some So we have to look downward.
We may use or if we have the constant We will see later, that this not interesting, see Definition 2.1 (page 2.1) and Remark 2.1 (page 2.1). This is a contradiction made of (graphically speaking) two paths of length one each. Interesting contradictions are made of slightly more complicated paths, like the Yablo Cell meaning
Basically, the only way to express that is impossible, is by the logical formula
A slightly different argument: How can we express that is wrong in a finite acyclic graph? Consider the set of minimal elements below , and the formula corresponding to the graph in DNF of these elements. This might be
(the graph is trivial, just
or a tautology or contradiction.
This argument is not totally satisfactory. For this reason, we reformulate:
We examine graphs, where there existe a node such that as well as -x will lead to a contradiction.
We want to make and contradictory. We know that this is impossible in finite graphs, so we have to show how we solve the problem in the (infinite) limit. We proceed inductively as Yablo did, but we also see that we basically have no other choice (modulo minor modifications).
So we want to express by an infinite graph that both and are inconsistent. As we saw in Section 1.3.1 (page 1.3.1), this must be expressed by suitable contradictions. Therefore, we treat such contradictions as basic building blocks, as Yablo did.
The problems are then
- (1)
This view is too abstract, we have to look at details, as they are given in the proof of validity for Yablo’s construction, or our reconstruction in Section 3 (page 3).
A summary: we have to look at instances.
- (2.1)
- (2.2)
- (2.4)
1.4 The “Right” Level of Abstraction
1.4.1 Introduction
The basic elements in Yablo’s construction are negative arrows, from which Yablo Cells are built. We show here that we can build arbitrarily complex structures equivalent to negative arrows, and as a matter of fact to any propositional logical operator. This suggests that the right level of abstraction to consider more general Yablo-like structures is not the level of single arrows, but rather of paths.
Remark 1.8
In the diagrams in Diagram 1.3 (page 1.3) the inner path is barred by , and the outer path contains 3 or 2 negations.
Definition 1.4
We define negations and their type.
A negation (diagram) is a diagram all of whose arrows are of the kind.
We define negation diagrams, or, simply, negations, and their type.
An arrow is a negation of type 0.
all negations inside are of type and at least one negation inside is of type
Remark 1.9
Thus, it seems very difficult to describe Yablo type diagrams on the level of single arrows. It seems to be the wrong level of abstraction.
We now give some examples.
1.4.2 Some Examples
Example 1.8
Negation (1)
so and
Negation (2)
so and continue as above.
Identity
so and
TRUE
Thus, and
so and
Diagram 1.3
Propositional Formulas 1
Diagram 1.4
Propositional Formulas 2
1.5 Prerequisites
We work with pure conjunctions of positive or negative propositional variables, like except for the following:
we may use a constant TRUE or
we may use formulas of the type
so the following are also permitted: etc.
This is used only to “neutralize” unwanted parts of the graph we characterize, see Fact 4.4 (page 4.4).
- (2)
- (3)
- (4)
2 Elementary Cells
2.1 Introduction
Elementary cells have to be simple constructions providing contradictions (in some cases, e.g. when the origin, usually called or so, cannot have a certain truth value, as this will lead, within the cell, to a contradiction).
Thus
They must be contradictory, and thus exclude at least one instance.
In addition, they should be simple, e.g. they will not contain a sequence or so without branching, as we could replace this with
(We will see later how to relinquish this requirement.)
More importantly, we must be able to construct a Yablo-like structure, built up by this type of elementary cell only (and perhaps some simpler auxiliary structures). They must serve as building blocks for a suitable stucture. This prevents “cheating”, letting other constructions do the crucial work.
- (4)
2.2 Basic Discussion
See Section 7.3 in [Sch22] for more material.
We try to describe here the basic constructions of contradictions.
As said in Fact 1.1 (page 1.1) (and re-written graph-theoretically), these consist of two branches with common origin, which meet again, and have different polarity. We call such constructions cells.
Second, we will define a simple hierarchy of such cells, and will allow a cell only the use of simpler cells as substructure. This prevents “cheating”. (See Section 2.1 (page 2.1).) Thus, we look at those cells only which allow to construct a Yablo-like structure without the use of more complicate cells.
Definition 2.1
We will argue that it suffices to consider these types of cells, see Remark 2.3 (page 2.3) and Section 2.4 (page 2.4). Furthermore, we can also neglect all but the (slightly generalized) Yablo Cells, as we will see in the present section.
Remark 2.1
An escape problem with for the trivial contradiction, so (See Definition 2.1 (page 2.1).) If we continue with so etc. we have an escape sequence etc., which never meets a contradiction.
Suppose we have an escape problem, like thus
This, however, is not possible.
Example 2.1
We consider now some simple, contradictory cells. They should not only be contradictory for the case but also be a potential start for the case without using more complex cells. (Otherwise, we postpone the solution, and may forget the overly simple start.)
We will consider in Section 2.5 (page 2.5) a more complicated contradiction cell (“Diamond”), and discuss here only simple cells.
Note that the following considerations apply also to cells formed by paths consisting of more than one arrow, and not only to cells formed by single arrows. See Section 2.4 (page 2.4). In particular, also in more complicated cells all sides have to negative.
The cell with 2 arrows.
It corresponds to the formula graphically, it has a positive and a negative arrow from to so exactly one of and is negative.
If is positive, we have a contradiction.
Of course, appending at a Yablo Cell (see below, case (2.3)) may be the beginning of a contradictory structure, but this is “cheating”, we use a more complex cell.
Cells with 3 arrows.
In the graph with not annotated arrows, they have the form -z, -z, where the lines may stand for or
Again, we want a contradiction for positive, so we need an at For the same reason, the number of negative arrows may not be even. So assume the number of negative arrows is one or three.
Three negative arrows:
This is the original type of contradiction in Yablo’s construction
This will be discussed in detail in the rest of the paper. But we see already that both paths, and change sign, so and will be positive, appending the same type of cell at and solves the problem (locally), and offers no escape.
We call sometimes the head of the Yablo cell, its foot, and its knee.
If we combine Yablo Cells, the knee for one cell may become the head for another Cell, etc.
One negative arrow:
We will show that all cases have an escape possibility, along the positive arrow (or arrows) originating in So they are not suitable for a construction using only this type of cell.
Consider the case
Upper part:
Lower part:
Consider the case
Consider the case
Here,
Diagram 2.1
Simple Contradictions
Diagram 2.2
Diagram ESC-1
Diagram 2.3
Diagram ESC-2
Diagram 2.4
Diagram ESC-3
Remark 2.2
The distinction between and i.e. between knee and foot, is conceptually important. In the case we have at a complete contradiction, at we have not yet constructed a contradiction. Thus, if we have at again an (as at this becomes an and we have to construct a contradiction for all (or not only for Otherwise, we have an escape possibility for Obviously, the contradiction need not be immediate at the important property is that ALL paths through all lead to a contradiction, and the simplest way is to have the contradiction immediately at - as in Yablo’s construction, and our saw blades.
See also the discussion of “Saw Blades” in [Sch22], Section 3.
As we work for a contradiction in the case x-, too, the simplest way to achieve this is to have a negative arrow and at again an This gives a chance to construct a contradiction at Of course, we have to construct a contradiction at too, as in the case we have an at
Of course, we may have branches originating at which all lead to contradictions in the case so (resp. at is possible, too.
But, for the simple construction, we need between and and at And this leads to the construction of contradictions for all (or as just mentioned above.
2.3 Values for y and z in Yablo Cells
We consider here what happens when we give and values In addition, we consider the variants and
Remark 2.3
False, True
We present a systematic treatment of variants of the “environment” of the Yablo triangle (and the diamond).
In the Yablo construction, we attach to the Yablo triangle to and constructions, which make We also consider here the cases where we end by or and set e.g. etc., see (3.1) below.
Note that will always hold.
We consider the following requirements:
Do we have a contradiction for i.e. will not be true in the Yablo Cell?
Will both branches lead to possible contradictions for
Do we obtain
- (2.4)
We consider the cases in all possible combinations:
Case is equivalent to case
Thus, we look at e.g. case i.e. and etc., and consider whether the requirements in (2) above are satisfied.
We first look only at the minimal requirements
has to be contradictory, requirement (2.1) above.
must not be classical (i.e. not or requirement (2.3) above.
The cases:
(2) (requirement (2.3))
Consider so for any so these cases are classical.
Consider For is classical, so is, too.
so
Thus, is the only non-classical case among
Consider In cases so too.
In case so
In case so
So is the only classical case among the
So the interesting cases left are
(1) (requirement (2.1))
We now check for local contradiction in the case i.e. must not be
may only be iff But if then only in case
Next, we have to look at and see if every path from leads to a contradiction, requirement (2.2) above. So neither nor must be
So and will not work.
Consider i.e.
Then in and so this does not work.
so this works.
so this works.
Consider i.e.
Then or and all cases work, except
In summary: in exactly the following cases: so they are excluded.
We summarize the results for the three requirements above:
(2.3): are ok.
(2.2): are not ok, so we learn nothing new beyond requirement (2.3).
(2.1): Only is eliminated, so we learn nothing new beyond requirement (2.3).
In summary, only satisfy all three requirements, (2.1), (2.2), (2.3).
We look at these cases.
Here, and in (3.6), stands for which is of type likewise as we will append in the full structure at and a construction with value
In all cases
Thus,
Thus,
Thus,
identical to
Thus,
Thus,
We look at the Diamond.
Thus,
The modified Diamond is the same, only so and so we have a Yablo cell.
Summary of the (in this context) important constructions
2.4 Yablo Cells Formed By Paths
We consider a triangle , , where etc. may be composite paths. We know from Section 2.2 (page 2.2) that the paths have to be negative, but they may be composed of positive and negative arrows.
We require that is of type and that the triangle is contradictory for the case as well as for both paths und have to lead to a contradiction (conditions (C1) and (C2)). Note that and have to be the first nodes on above paths to be contradictory for - otherwise we will not reach (or the path is barred before.
2.4.1 Branching points
Branching at
Suppose we branch at so we have, in addition to the triangle, some
We know that is negative, so, if then This is the situation of the Yablo construction, see Section 3 (page 3). So just a short comment here.
This however, is possible only a finite number of times, e.g. then etc., see Example 1.6 (page 1.6). This constructs an infinite sequence of choices which offers an escape possibility.
Thus we need infinite branching (and non-classical logic).
2.4.2 Paths
Additional branching on , e.g. , ,
Case 1: OR
This generates a copy of the structure for
We construct for as for , i.e.
Case 2: AND
So OR, so we make a copy of the structure for consider e.g. , ,
Note: We simplify, suppose e.g. that we have then we do not need another contradiction at the one at suffices (condition (C1)). We need, however, that all branches lead to contradictions (condition (C2)).
We branch on , and have , .
Case 1: OR:
We make a copy of the structure for
Case 2: AND:
So x- OR:
We make a copy of the structure for
Consider e.g. , , .
Diagram 2.5
2.5 The Problem with Diamonds
Remark 2.4
The simplified version with recursion on one side only, here on the left only, is nothing but the Yablo triangle:
Diagram 2.6
Nested Diamonds, Details
Diagram 2.7
3 Combining Cells
3.1 Overview
We follow Yablo’s construction.
(C1): We have to make contradictory.
(This is not an important simplification, as more complicated constructions will finally use our simple one, too. See e.g. Diagram 1.2 (page 1.2) and its discussion.)
As Yablo did, we set and
(C2) We have to make contradictory.
Probabilistically, of the possibilities at are open.
Etc.
See below and Section 3.3.3. in [Sch23b].
3.2 In More Detail
We now consider the inductive construction of the Yablo structure, see Diagram 3.1 (page 3.1) and Diagram 3.2 (page 3.2)
abbreviated
has to lead to a contradiction for
preparation,
finish,
(branch has to be contradicted for add
has to lead to a contradiction for
preparation,
finish,
(branch has to be contradicted for add
(branch has to be contradicted for add
Construction 3.1
Summary:
Every is head of a Yablo cell with knee
The construction has infinite depth and branching.
Remark 3.1
Our construction is somewhat arbitrary. We construct a contradiction for by and use the arrow as part of the contradiction for This, of course, is a special case.
Consider Diagram 3.3 (page 3.3). Here, the contradiction for is as above, but we begin a totally new contradiction for likewise for etc.
But, we see, the old construction of Diagram 3.1 (page 3.1) and Diagram 3.2 (page 3.2) is part of the more general construction of Diagram 3.3 (page 3.3), so the old construction is really minimal.
We have to be a bit careful here. If we set etc. then we have etc. Instead, we may set so we have as desired, or we append at recursively the same construction - as in “Saw Blades”.
Diagram 3.1
Diagram Inductive Construction 1
Diagram 3.2
Diagram Inductive Construction 2
Diagram 3.3
Diagram Inductive Construction Core
4 Result